OpenStudy (anonymous):

Let f(x)=x-2 and g(x)=x2(squared)-7x-9. Find f(g(-1)).

OpenStudy (anonymous):

g(-1) = +1 f(1) = 1 - 2 = -1

OpenStudy (anonymous):

answer choices are: -21 -3 3 21

OpenStudy (anonymous):

@ginochen

OpenStudy (anonymous):

sorry, g(-1) = (-1)(-1) -7(-1)-9 = 1+7-9=-1 f(-1)=-1-2=-3